The accumulated amount after eight years of monthly payments of R1 900 each into an account earning 9,7% interest per year, compounded monthly, is

[1] R274 069,25.

[2] R182 400,00.

[3] R126 532,64.

[4] R395 077,74.

[5] none of the above. 


1
Expert's answer
2021-05-31T15:00:50-0400

Periodic payment PMT=1900

Rate of interest per period "i= \frac{0.097}{12}= 0.0080833"

Number of periods "n = 12 \times 8 = 96"

By future value formula

"FV = PMT(\frac{(1+i)^n-1}{i})\\

FV = 1900(\frac{(1+0.0080833)^{96}-1}{0.0080833} )\\

FV = 1900 \times 144.2467 \\

FV = 274068.76"

Answer: [1] R274 069,25.